OpenStudy (anonymous):

two angles are complementary that measure 90 degrees. the measure of one angle is 1 1/2 times the measure of the other. Find the mearsure of each angle.

OpenStudy (anonymous):

36 and 54

OpenStudy (anonymous):

let the angles be x and 1 1/2x \[x+1\frac{1}{2}x=90\] \[2\frac{1}{2}x=90\] \[\frac{5}{2}x=90\] \[x=\frac{90*2}{5}\] \[x=36\] angles are 36 and (3/2)*36=54